OpenStudy (anonymous):

Find all solutions in the interval [0,2) Sec^2x-2=tan^2x Is the answer no solution?

OpenStudy (anonymous):

sec^2x = tan^2x + 1, so tan^2x + 1 - 2 = tan^2x -1 = 0 (false statement) there are no solutions

OpenStudy (anonymous):

Awww yess lawddd! Thank you man :)